40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 25 15:33:55 2012, janosch, Update, LSC, ringdown measurement Rindown1.png
    Reply  Wed Jul 25 16:31:01 2012, Jenne, Update, LSC, Yarm green locking to arm - PDH box saturating 
    Reply  Wed Jul 25 22:16:36 2012, Koji, Update, LSC, ringdown measurement 
Message ID: 7029     Entry time: Wed Jul 25 15:33:55 2012     Reply to this: 7030   7033
Author: janosch 
Type: Update 
Category: LSC 
Subject: ringdown measurement 

We did our first ringdown measurement on the Y arm. First we tried to keep the arm locked in green during the ringdown, but for some reason it was not possible to get the cavity locked in green. So we decided to do the first measurement with infrared locked only.

For the measurement we had to change the LSC model to acquire the C1:LSC-TRY_OUT_DQ at higher sampling frequency. We changed the sampling frequencies of C1:LSC-{TRX,TRY}_OUT_DQ from 2048Hz to 16384Hz.

The measurement was done at GPS 1027289507. The ringdown curve looks very clean, but there seem to be two time constants involved. The first half of the curve is influenced by the shutter speed, then curvature is changing sign and the ringdown is likely taking over. We will try to fit a curve to the ringdown part, but it would certainly be better to have a faster shutter and record a more complete ringdown.

Rindown1.png

 

ELOG V3.1.3-